# Defect design in ferroelectrics – new insights on agglomeration

### v3

@@ Abstract
Functional properties of ferroelectrics and their change with time depend crucially on the defect structure. In particular, point defects and bias fields induced by defect dipoles modify the field hysteresis and play an important role in fatigue and aging. However, a full understanding on how order, agglomeration and strength of defect dipoles affect phase stability and functional properties is still lacking. To close these gaps in knowledge, we screen these parameters by ab initio based molecular dynamics simulations with the effective Hamiltonian method for the prototypical ferroelectric material (Ba,Sr)TiO3. Our findings suggest that the active surface area of the defects, rather than the defect concentration is the decisive factor. For a fixed defect concentration, clustering reduces the active surface area and thus the defect-induced changes of phase stability and field hysteresis. Particularly planar agglomerates of defects appear as promising route for the material design as their impact on the field hysteresis can be controlled by the field direction and their impact on the phase stability shows a cross-over with the strength of the defect dipoles. For this agglomeration, we furthermore show that a pinched double-loop field hysteresis allows for large recoverable stored energy which can outperform the response of pristine (Ba,Sr)TiO3 even in its paraelectric phase.
@@ 2.1 Molecular dynamics simulations
In (Ba,Sr)TiO3, the transition temperatures and the polarization in all ferroelectric phases are reduced with the Sr concentration, see Fig. 12. Without loss of generality, we focus on Ba0.7Sr0.3TiO3 with its Curie temperature close to room temperature using a random distribution of Sr in combination with the chemical modulation $V^{\text{mod,inho}}$ which depends on the Sr concentration in the unit cell and a pressure correction of $-1.2$ GPa, cf. Ref. 36.
The other terms in Eq. (1) are the self-energy, $V^{\text{self}}$, the long-range dipole-dipole interaction, $V^{\text{dpl}}$, the short-range harmonic interaction between local modes, $V^{\text{short}}$, the elastic energy from homogeneous and inhomogeneous strains, $V^{\text{elas,homo}}$ and $V^{\text{elas,inho}}$, the coupling terms between local soft mode and strain, $V^{\text{coup,inho}}$ and $V^{\text{coup,inho}}$. As the strain is optimized internally, only the dynamics of the dipole moments are explicitly treated in molecular dynamics simulations, using $M^{*}_{\text{dpl}}$ the effective mass of the local soft mode. This coarse-graining is illustrated in Fig. 1.
@@
A supercell with $L_{x}\times L_{y}\times L_{z}=48\times 48\times 48$ u.c. is combined with periodic boundary conditions and the dipoles are initialized with a normal distribution with ${\left\langle{u_{x}}\right\rangle}={\left\langle{u_{y}}\right\rangle}={\left\langle{u_{z}}\right\rangle}=0.00$ and ${\left\langle{u_{i}^{2}}\right\rangle}-{\left\langle{u_{i}}\right\rangle}^{2}=0.03$2 at 350 K or with ${\left\langle{u_{x}}\right\rangle}={\left\langle{u_{y}}\right\rangle}={\left\langle{u_{z}}\right\rangle}=0.10$ and ${\left\langle{u_{i}^{2}}\right\rangle}-{\left\langle{u_{i}}\right\rangle}^{2}=0.01$2 at 50 K to perform cooling and heating simulations. For each temperature, we use a time step of $\Delta t=$2 fs and thermalize the system for 40000 steps in the canonical ensemble using the Nosé-Poincaré thermostat before we calculate the average properties in additional 20000 steps. The final configuration is the input for the next temperature step using $\Delta T=\pm 1$ K. We monitor the absolute values of the macroscopic polarization along direction $i$ ($|{\left\langle{P_{i}}\right\rangle}|$), and the mean of the magnitude of the local polarization ${\left\langle{|P_{i}|}\right\rangle}$. Both quantities differ in the presence of domains. Unlike stated otherwise, the temperatures of ferroelectric to ferroelectric transitions are identified from the maximum/minimum of strain in the high/low-temperature phase. The paraelectric-ferroelectric transition temperature is determined from the maximal derivative of polarization with respect to temperature.
@@ 2.2 Defect dipoles
Distribution Concentration (%) Active surface area
@@ 3.1 Phase diagrams
Figure 5: Schematic illustration of the active surface areas of a stripe of non-polar defect dipoles (black) of width $w$, height $n_{z}$ and spanning the whole space along $y$ (not shown). Green and gray lines indicate the active surface areas for the polarization components $P_{z}$ and $P_{x}$ of the surrounding free dipoles. (a) In the T phase, the polarization aligns along $y$ to reduce the active surface area ($w>n_{z}>0$). (c) In the R phase, the polarization component $P_{z}$ faces a large active area and can lower the depolarization field by the formation of domains. (b) Changes of total energy by 1% non-polar defects ($|u_{D}|=0.0$, blue) and polar defects ($|u_{D}|=0.1$ $\hat{z}$, red) in the T phase (250 K). Mean values of 20000 time steps and their standard deviations are given.
@@
While the main trends of the impact of defects on the system are thus related to the active surface areas of the individual defects, this concept is only exact for well-separated defects. In addition, also the distances and dilution of defects impact phase stability and dipole configuration. Defects randomly distributed in 3D mainly reduce the dipole moments of their nearest neighbors along the polarization direction, e.g. about 4.4% dipoles are reduced to about 7.2 $\mu C/cm^{2}$ in the presence of 2% defect, cf. Fig. 18. As the defects start to cluster in 2D, more and more dipoles are reduced and the coupling between defects and dipoles becomes more long-ranged. Already a stripe with $w=6$ u.c. (i.e. 0.5% defects) induces gradual rotations of dipoles in at least 3 layers next to the agglomerate. If a second defect cluster is within this distance, no bulk-like macroscopic polarization can form in the gap between them and both clusters act as one cluster with a joint effective active surface area, cf. Fig. 17. Above a critical width of about $w=12$ u.c., negative domains with a diameter of $8$ u.c. form at the defect agglomerate, and for $w=24$ u.c. the dipoles normal to the defect are modified in the whole system by the formation of domains with R71 walls along $[110]$ due to depolarization fields [2], cf. Fig. 21.
@@
Figure 6: Temperature - defect concentration phase diagrams for polar defects ($|u_{D}|=$0.1 Å$\hat{z}$). Symbols mark transition temperatures for cooling (blue) and heating (red). Note that no C-T transition can be detected for large defect concentrations and magnitudes beyond the critical point. Random ordering (solid lines) and agglomeration (dashed lines) are compared for (a) defects in 3D and (b) defects in a $z$-plane. Numbers give the approximate number of active surface units in all phases for (a) cube and random distribution (italic) and (b) for stripes.
@@
In case of 2D agglomeration, the effective active surface area along $x$ and $y$ is again zero and the transition temperatures to O or R phases under cooling are not sensitive to the defect concentration. On the other hand, randomly distributed defects have an active surface area at their non-polar interfaces and thus induce a reduction of $T_{C}^{\text{T-O}}$ of more than $55$ K for 2 % defects and also $T_{C}^{\text{O-R}}$ goes down with the defect concentration, see Fig. 6 (a). For 2D agglomerates with, e.g., $|u_{D}|=0.1$ (Å) $\hat{x}$, the defects act similar to non-polar defects, cf. Fig. 14, with polarization along $x$ and $xy$ in the T and O phases and small changes of the transition temperatures, while $T_{C}^{\text{O-R}}$ decreases by $36$ K ($1$ of defects) and by $98$ K for a full defect plane for the R phase with polarization along $z$. Comparison between these defects and the non-polar ones shows that the former induces a slight increase of $T_{C}^{\text{C-T}}$ and a reduction of the thermal hysteresis at the T-O transition, while only in the latter case $T_{C}^{\text{T-O}}$ increases slightly for large defect concentrations. All these observations can be related to the small internal fields along $x$.
@@
Figure 7: Impact of the strength of 1 % defects ($|u_{D}|$ $\hat{z}$) on the ferroelectric phase diagram. Colors highlight transition temperatures for cooling (blue) and heating (red). Agglomeration (solid lines) and random ordering (dashed lines) are compared for (a) defects in a cube or 3D and (b) defects in one $z$-plane. For both the T and O phases in (b), horizontal ($=$) and vertical (||||| |) hatches mark $P_{z}=0$ and $|P_{z}|>0$, respectively. (c) and (d) show corresponding time-averaged local dipoles at 200 K for weak defect dipoles ($|u_{D}|=0.03$ $\hat{z}$) and strong defect dipoles ($|u_{D}|=0.1$ $\hat{z}$), respectively. Note that beyond $0.04$ the sample with 2D random defects is beyond the critical point of the C-T transition.
@@
For 2D agglomerates and defects randomly distributed in plane with polarization along $z$, where is a cross-over from weak to strong defect dipoles, see Fig. 7 (b). Below $u_{z}=0.03$, e.g. about 27% of the saturation polarization of the pristine material, the defect dipoles act similar to non-polar defects with polarization in the T and O phases pointing along $P_{y}$ and $P_{y}=P_{x}$, no effective active surface area and no change of $T_{C}^{\text{C-T}}$ and $T_{C}^{\text{T-O}}$. Only in the R phase, one polarization component points parallel to the defects and $T_{C}^{\text{O-R}}$ increases with their strength. Thereby the O–R transition becomes continuous, cf. Fig.22 (a)–(b). These weak defects are smaller than the surrounding free dipoles above $T_{C}^{\text{C-T}}$, thus inducing depolarization fields. For strong defects ($u>0.03$), the polarization in the T phase point along $z$ and $T_{C}^{\text{C-T}}$ increases with defect strength up to the critical point, while $T_{C}^{\text{O-R}}$ and $T_{C}^{\text{T-O}}$ decrease with the internal defect field. As soon as their magnitude is larger than the typical values of free dipoles slightly above $T_{C}^{\text{C-T}}$, these defects start to act as an internal electric field, promoting the formation of the ferroelectric phase along that direction, cf. discussion in D.
@@ 3.2 P-E Hystereses
Figure 8: Changes of the field hysteresis at $T=250$ K if going from the pristine material (red) to systems with different distribution of 1 % defects (green: cube; blue and cyan: stripes and randomly in $\hat{z}$-planes; purple: randomly) for (a) non-polar defects and (b)–(c) polar defect dipoles ($|u_{D}|=$0.1 Å$\hat{z}$) for the electric field (b) perpendicular ($E_{x}$) and (c) collinear ($E_{z}$) to the defect dipoles. Only the polarization component $P_{i}$ along the field direction is shown.
How do the different defect configurations affect the field response of the material? As reference, red lines in Fig. 8 show the field hysteresis of the pristine material in the T phase at 250 K. The $P-E$ curve is symmetric, centered at the origin, and has large coercive fields and remanent polarization. With temperature, hysteresis and polarization are reduced, and finally in the paraelectric phase, the field response is linear, see Fig. 13 for 350 K. The characteristics in the ferroelectric phase change considerably with the strength and ordering of defects, see other colors in Fig. 8 and the corresponding strain–field curves in Fig. 23. Non-polar or weak defect dipoles, reduce remanent polarization ($|P_{r}^{\pm}|$) and particularly the width of the hysteresis loops ($\Delta E_{c}$) for all orderings while the hysteresis stays symmetric as shown in subfigure (a). Analogous to the phase diagrams, the differences between the configurations are related to the active surface area of the defects. The active surface area and the impact on the field hysteresis is small for cubic agglomerates and 2D agglomerates with the field applied in the defect plane. In these cases the changes of $\Delta E_{c}$ (and $P_{r}$) are below $18$ ($3$) and $20$ ($3$), respectively.
Larger modifications are induced by randomly ordered defects that have a larger active surface area. The largest changes are induced by 2D agglomerates if the field is applied along their normal, i.e.  along $z$. For this geometry, the depolarization field induced by the defects acts as restoring force on the polarization direction. Therefore, the polarization switches reversibly between the tetragonal phase in plane with $P_{x}$ and a monoclinic state with ±$\pm P_{z}>P_{x}$, cf. Fig. 24. The $P_{z}(E_{z})$ hysteresis is pinched to a double-loop. Thereby, for the chosen defect concentration of 1%, the randomly distributed defects in one plane impact the surrounding dipoles more than stripes, cf. discussion on Fig. 18, and stabilize a reversible linear change of $P_{z}$ with the field for more than ±$\pm 96$ kV/cm.
@@
Figure 9: Characteristics of the field hysteresis for 1 % (a,c,e) non-polar and (b,d,f) polar defects ($|u_{D}|=$0.1 Å$\hat{z}$) at $T=250$ K for $E=150$ kV/cm sorted by efficiency. In (a)–(b) $W_{\text{rec}}$ (green) and $W_{\text{loss}}$ (gray hatched) and the efficiency $\eta$ (black dots, right axes), in (c)–(d) the remanent polarization $|P_{r}|$ and in (e)–(f) the width of the hysteresis ($\Delta E_{c})$ and the bias field $E_{\text{bias}}$ for different defect distributions are compared to the pristine material. Red and blue correspond to positive and negative field directions which differ for the collinear alignment of defects and the external field ($E_{z}$).
Figure 9 compares the characteristics of the hysteresis loops for all defect configurations at 250 K. At this temperature, the pristine material is in the T phase and a work loss of about 3 J/cm-3 is dissipated if the polarization switches from $P_{r}^{-}$ to the saturation polarization $P_{\text{max}}^{x}$ and vice versa for negative field directions, while the recoverable work as well as reversible changes of polarization and strain without switching are negligible, in qualitative agreement with previous reports [50]. On the one hand, non-polar or weak defects agglomerated in cubes or 2D with parallel applied fields induce a small reduction of the loss in the ferroelectric phases only. The efficiency $\eta$ is thus approximately constant between these configurations and increases slightly for random defects and 2D defects in perpendicular applied fields. On the other hand, large efficiencies are possible for the double loop hysteresis found for 2D ordering. Thereby, the smallest hysteresis in case of defects randomly distributed in one plane allows for an efficiency of about 94 %. Analogously, pinched hysteresis with reduced losses and maximal $\Delta P$ and $\eta$ are stabilized by strong defects along $z$ in a field along $x$ as long as the active surface area is not reduced by clustering in 3D. Maximal recoverable energy and efficiency of 1.07 J/cm-3 and 96 % are possible for defects randomly distributed in the sample. Finally, for strong defects and fields along $z$, one has to distinguish between positive and negative field directions. In the former case shown in red, no polarization switching can be induced during unipolar field loops for randomly distributed defects and thus the used definition of loss and efficiency would predict optimal configurations. However, this is not relevant as the possible recoverable energy as well as changes of $P$ and strain are negligible. In case of the negative field shown in blue, maximal values of $W_{\text{rec}}$ and a moderate efficiency are possible due to the shifted and reduced hysteresis for random defect distributions.
@@
It is important that these large changes of the field hysteresis by defects are expected in the full temperature range of the ferroelectric phase. Well above $T_{C}^{T-C}$, in the paraelectric phase, losses are also absent, but at the same time the saturation polarization is small. Therefore, $W_{\text{rec}}$ (cf. Eqn (2)) is only 0.24 J/cm3 (0.75 J/cm3) for $E_{\text{max}}=50$ kV/cm ($E_{\text{max}}=100$ kV/cm) at 350 K. These values of $W_{\text{rec}}$ can be outperformed in the presence of 1% polar defects ($|u_{D}|=0.1$ $\hat{z}$) randomly distributed in 2D with perpendicular applied fields $E_{x}$ (0.42 and 0.77 J/cm3 for $E_{max}=50$ and $100$ kV/cm, respectively), see Fig. 13.

### v4

# Point defect design in (Ba,Sr)TiO3 – new insights on agglomeration

@@ Abstract
Functional properties of ferroelectrics and their changes with time depend crucially on the defect structure. In particular, point defects and bias fields induced by defect dipoles modify the field hysteresis and play an important role in fatigue and aging. However, a full understanding on how order, agglomeration and strength of defect dipoles affect phase stability and functional properties is still lacking. To close these gaps in knowledge, we screen these parameters by ab initio based molecular dynamics simulations with the effective Hamiltonian method for the prototypical ferroelectric material (Ba,Sr)TiO3. Our findings suggest that the active surface area of the defects, rather than the defect concentration is the decisive factor. For a fixed defect concentration, clustering reduces the active surface area and thus the defect-induced changes of phase stability and field hysteresis. Particularly planar agglomerates of defects appear as promising route for the material design as their impact on the field hysteresis can be controlled by the field direction and as their impact on the phase stability shows a cross-over with the strength of the defect dipoles. For this agglomeration, we show that the recoverable stored energy can outperform the response of pristine (Ba,Sr)TiO3 even in its paraelectric phase due to a pinched double-loop field hysteresis.
@@ 2.1 Molecular dynamics simulations
In (Ba,Sr)TiO3, the transition temperatures and the polarization in all ferroelectric phases are reduced with the Sr concentration, see Fig. 12. Without loss of generality, we focus on Ba0.7Sr0.3TiO3 with its Curie temperature close to room temperature using a random distribution of Sr in combination with the chemical modulation $V^{\text{mod,inho}}$ which depends on the Sr concentration in the unit cell and a pressure correction of 1.2-1.2 GPa, cf. Ref. 36.
The other terms in Eq. (1) are the self-energy, $V^{\text{self}}$, the long-range dipole-dipole interaction, $V^{\text{dpl}}$, the short-range harmonic interaction between local modes, $V^{\text{short}}$, the elastic energy from homogeneous and inhomogeneous strains, $V^{\text{elas,homo}}$ and $V^{\text{elas,inho}}$, the coupling terms between local soft mode and strain, $V^{\text{coup,homo}}$ and $V^{\text{coup,inho}}$. As the strain is optimized internally, only the dynamics of the dipole moments are explicitly treated in molecular dynamics simulations, using $M^{*}_{\text{dpl}}$ the effective mass of the local soft mode. This coarse-graining is illustrated in Fig. 1.
@@
A supercell with $L_{x}\times L_{y}\times L_{z}=48\times 48\times 48$ u.c. is combined with periodic boundary conditions and the dipoles are initialized with a normal distribution with ${\left\langle{u_{x}}\right\rangle}={\left\langle{u_{y}}\right\rangle}={\left\langle{u_{z}}\right\rangle}=0.00$ Å and ${\left\langle{u_{i}^{2}}\right\rangle}-{\left\langle{u_{i}}\right\rangle}^{2}=0.03$ Å2 at 350 K or with ${\left\langle{u_{x}}\right\rangle}={\left\langle{u_{y}}\right\rangle}={\left\langle{u_{z}}\right\rangle}=0.10$ Å and ${\left\langle{u_{i}^{2}}\right\rangle}-{\left\langle{u_{i}}\right\rangle}^{2}=0.01$ Å2 at 50 K to perform cooling and heating simulations. For each temperature, we use a time step of $\Delta t=$2 fs and thermalize the system for 40000 steps in the canonical ensemble using the Nosé-Poincaré thermostat before we calculate the average properties in additional 20000 steps. The final configuration is the input for the next temperature step using $\Delta T=\pm 1$ K. We monitor the absolute values of the macroscopic polarization along direction $i$ ($|{\left\langle{P_{i}}\right\rangle}|$), and the mean of the magnitude of the local polarization $\left\langle{|P_{i}|}\right\rangle$}. Both quantities differ in the presence of domains. Unlike stated otherwise, the temperatures of ferroelectric to ferroelectric transitions are identified from the maximum/minimum of strain in the high/low-temperature phase. The paraelectric-ferroelectric transition temperature is determined from the maximal derivative of polarization with respect to temperature.
@@ 2.2 Defect dipoles
Distribution
Concentration
Active surface area
@@ 3.1 Phase diagrams
Figure 5: Schematic illustration of the active surface areas of a stripe of non-polar defect dipoles (black) of width $w$, height $n_{z}$ and spanning the whole space along $y$ (not shown). Green and gray lines indicate the active surface areas for the polarization components $P_{z}$ and $P_{x}$ of the surrounding free dipoles. (a) In the T phase, the polarization aligns along $y$ to reduce the active surface area ($w>n_{z}>0$). (c) In the R phase, the polarization component $P_{z}$ faces a large active area and can lower the depolarization field by the formation of domains. (b) Changes of total energy by 1% non-polar defects ($|u_{D}|=0.0$ Å, blue) and polar defects ($|u_{D}|=0.1$ Å $\hat{z}$, red) in the T phase (250 K). Mean values of 20000 time steps and their standard deviations are given.
@@
[…]
@@
Figure 6: Temperature - defect concentration phase diagrams for polar defects ($|u_{D}|=$0.1 $\hat{z}$). Symbols mark transition temperatures for cooling (blue) and heating (red). Note that no C-T transition can be detected for large defect concentrations and magnitudes beyond the critical point. Random ordering (solid lines) and agglomeration (dashed lines) are compared for (a) defects in 3D and (b) defects in a $z$-plane. Numbers give the approximate number of active surface units in all phases for (a) cube and random distribution (italic) and (b) for stripes.
@@
In case of 2D agglomeration, the effective active surface area along $x$ and $y$ is again zero and the transition temperatures to O or R phases under cooling are not sensitive to the defect concentration. On the other hand, randomly distributed defects have an active surface area at their non-polar interfaces and thus induce a reduction of $T_{C}^{\text{T-O}}$ of more than $55$ K for 2 % defects and also $T_{C}^{\text{O-R}}$ goes down with the defect concentration, see Fig. 6 (a). For 2D agglomerates with, e.g., $|u_{D}|=0.1$ (Å) $\hat{x}$, the defects act similar to non-polar defects, cf. Fig. 14, with polarization along $x$ and $xy$ in the T and O phases and small changes of the transition temperatures, while $T_{C}^{\text{O-R}}$ decreases by $36$ K ($1$ % of defects) and by $98$ K for a full defect plane for the R phase with polarization along $z$. Comparison between these defects and the non-polar ones shows that the former induces a slight increase of $T_{C}^{\text{C-T}}$ and a reduction of the thermal hysteresis at the T-O transition, while only in the latter case $T_{C}^{\text{T-O}}$ increases slightly for large defect concentrations. All these observations can be related to the small internal fields along $x$.
@@
Figure 7: Impact of the strength of 1 % defects ($|u_{D}|$ $\hat{z}$) on the ferroelectric phase diagram. Colors highlight transition temperatures for cooling (blue) and heating (red). Agglomeration (solid lines) and random ordering (dashed lines) are compared for (a) defects in a cube or 3D and (b) defects in one $z$-plane. For both the T and O phases in (b), horizontal (==) and vertical ($||$) hatches mark $P_{z}=0$ and $|P_{z}|>0$, respectively. (c) and (d) show corresponding time-averaged local dipoles at 200 K for weak defect dipoles ($|u_{D}|=0.03$ Å $\hat{z}$) and strong defect dipoles ($|u_{D}|=0.1$ Å $\hat{z}$), respectively. Note that beyond $0.04$ Å the sample with 2D random defects is beyond the critical point of the C-T transition.
@@
For 2D agglomerates and defects randomly distributed in plane with polarization along $z$, where is a cross-over from weak to strong defect dipoles, see Fig. 7 (b). Below $u_{z}=0.03$ Å, e.g. about 27% of the saturation polarization of the pristine material, the defect dipoles act similar to non-polar defects with polarization in the T and O phases pointing along $P_{y}$ and $P_{y}=P_{x}$, no effective active surface area and no change of $T_{C}^{\text{C-T}}$ and $T_{C}^{\text{T-O}}$. Only in the R phase, one polarization component points parallel to the defects and $T_{C}^{\text{O-R}}$ increases with their strength. Thereby the O–R transition becomes continuous, cf. Fig.22 (a)–(b). These weak defects are smaller than the surrounding free dipoles above $T_{C}^{\text{C-T}}$, thus inducing depolarization fields. For strong defects ($u>0.03$ Å), the polarization in the T phase point along $z$ and $T_{C}^{\text{C-T}}$ increases with defect strength up to the critical point, while $T_{C}^{\text{O-R}}$ and $T_{C}^{\text{T-O}}$ decrease with the internal defect field. As soon as their magnitude is larger than the typical values of free dipoles slightly above $T_{C}^{\text{C-T}}$, these defects start to act as an internal electric field, promoting the formation of the ferroelectric phase along that direction, cf. discussion in D.
@@ 3.2 P-E Hystereses
Figure 8: Changes of the field hysteresis at $T=250$ K if going from the pristine material (red) to systems with different distribution of 1 % defects (green: cube; blue and cyan: stripes and randomly in $\hat{z}-$planes; purple: randomly) for (a) non-polar defects and (b)–(c) polar defect dipoles ($|u_{D}|=$0.1 $\hat{z}$) for the electric field (b) perpendicular ($E_{x}$) and (c) collinear ($E_{z}$) to the defect dipoles. Only the polarization component $P_{i}$ along the field direction is shown.
How do the different defect configurations affect the field response of the material? As reference, red lines in Fig. 8 show the field hysteresis of the pristine material in the T phase at 250 K. The PEP-E curve is symmetric, centered at the origin, and has large coercive fields and remanent polarization. With temperature, hysteresis and polarization are reduced, and finally in the paraelectric phase, the field response is linear, see Fig. 13 for 350 K. The characteristics in the ferroelectric phase change considerably with the strength and ordering of defects, see other colors in Fig. 8 and the corresponding strain–field curves in Fig. 23. Non-polar or weak defect dipoles, reduce remanent polarization ($|P_{r}^{\pm}|$) and particularly the width of the hysteresis loops ($\Delta E_{c}$) for all orderings while the hysteresis stays symmetric as shown in subfigure (a). Analogous to the phase diagrams, the differences between the configurations are related to the active surface area of the defects. The active surface area and the impact on the field hysteresis is small for cubic agglomerates and 2D agglomerates with the field applied in the defect plane. In these cases the changes of $\Delta E_{c}$ (and $P_{r}$) are below $18$ % ($3$ %) and $20$ % ($3$ %), respectively.
Larger modifications are induced by randomly ordered defects that have a larger active surface area. The largest changes are induced by 2D agglomerates if the field is applied along their normal, i.e.  along $z$. For this geometry, the depolarization field induced by the defects acts as restoring force on the polarization direction. Therefore, the polarization switches reversibly between the tetragonal phase in plane with $P_{x}$ and a monoclinic state with $\pm P_{z}>P_{x}$, cf. Fig. 24. The $P_{z}(E_{z})$ hysteresis is pinched to a double-loop. Thereby, for the chosen defect concentration of 1%, the randomly distributed defects in one plane impact the surrounding dipoles more than stripes, cf. discussion on Fig. 18, and stabilize a reversible linear change of $P_{z}$ with the field for more than $\pm 96$ kV/cm.
@@
Figure 9: Characteristics of the field hysteresis for 1 % (a,c,e) non-polar and (b,d,f) polar defects ($|u_{D}|=$0.1 $\hat{z}$) at $T=250$ K for $E=150$ kV/cm sorted by efficiency. In (a)–(b) $W_{\text{rec}}$ (green) and $W_{\text{loss}}$ (gray hatched) and the efficiency $\eta$ (black dots, right axes), in (c)–(d) the remanent polarization $|P_{r}|$ and in (e)–(f) the width of the hysteresis ($\Delta E_{c})$ and the bias field $E_{\text{bias}}$ for different defect distributions are compared to the pristine material. Red and blue correspond to positive and negative field directions which differ for the collinear alignment of defects and the external field ($E_{z}$).
Figure 9 compares the characteristics of the hysteresis loops for all defect configurations at 250 K. At this temperature, the pristine material is in the T phase and a work loss of about 3 J/cm3 is dissipated if the polarization switches from $P_{r}^{-}$ to the saturation polarization $P_{\text{max}}^{x}$ and vice versa for negative field directions, while the recoverable work as well as reversible changes of polarization and strain without switching are negligible, in qualitative agreement with previous reports [50]. On the one hand, non-polar or weak defects agglomerated in cubes or 2D with parallel applied fields induce a small reduction of the loss in the ferroelectric phases only. The efficiency $\eta$ is thus approximately constant between these configurations and increases slightly for random defects and 2D defects in perpendicular applied fields. On the other hand, large efficiencies are possible for the double loop hysteresis found for 2D ordering. Thereby, the smallest hysteresis in case of defects randomly distributed in one plane allows for an efficiency of about 94 %. Analogously, pinched hysteresis with reduced losses and maximal $\Delta P$ and $\eta$ are stabilized by strong defects along $z$ in a field along $x$ as long as the active surface area is not reduced by clustering in 3D. Maximal recoverable energy and efficiency of 1.07 J/cm3 and 96 % are possible for defects randomly distributed in the sample. Finally, for strong defects and fields along $z$, one has to distinguish between positive and negative field directions. In the former case shown in red, no polarization switching can be induced during unipolar field loops for randomly distributed defects and thus the used definition of loss and efficiency would predict optimal configurations. However, this is not relevant as the possible recoverable energy as well as changes of $P$ and strain are negligible. In case of the negative field shown in blue, maximal values of $W_{\text{rec}}$ and a moderate efficiency are possible due to the shifted and reduced hysteresis for random defect distributions.
@@
It is important that these large changes of the field hysteresis by defects are expected in the full temperature range of the ferroelectric phase. Well above $T_{C}^{T-C}$, in the paraelectric phase, losses are also absent, but at the same time the saturation polarization is small. Therefore, $W_{\text{rec}}$ (cf. Eqn (2)) is only 0.24 J/cm3 (0.75 J/cm3) for $E_{\text{max}}=50$ kV/cm ($E_{\text{max}}=100$ kV/cm) at 350 K. These values of $W_{\text{rec}}$ can be outperformed in the presence of 1% polar defects ($|u_{D}|=0.1$ Å $\hat{z}$) randomly distributed in 2D with perpendicular applied fields $E_{x}$ (0.42 and 0.77 J/cm3 for $E_{max}=50$ and $100$ kV/cm, respectively), see Fig. 13.